# Energy-Aware Routing to Large Reasoning Models

### v1

@@ III-B1 Battery random walk
$B_{t}^{\mathrm{my}}\;=\;\sum_{u=1}^{t-1}\bigl(R_{u}-C_{u}^{\mathrm{my}}\bigr)$
@@
$=t\,\overline{K}\,\operatorname{\mathbb{E}}_{X\sim f_{X}}\!\left[E_{\mathrm{LB}}(X)^{2}\right].$ (24)
Thus, the lumped-time myopic battery process is described by the mean, $\mu_{B_{\mathrm{my}}}$, and variance, $\sigma_{B_{\mathrm{my}}}^{2}$:
$\mu_{B_{\mathrm{my}}}$ $\coloneqq\operatorname{\mathbb{E}}[B_{t}^{\mathrm{my}}]/t=\overline{R}-\overline{C}_{\mathrm{LB}}$
$\sigma_{B_{\mathrm{my}}}^{2}$ $\coloneqq\operatorname{Var}(B_{t}^{\mathrm{my}})/t=\sigma_{R}^{2}+\overline{K}\operatorname{\mathbb{E}}_{X\sim f_{X}}\!\left[E_{\mathrm{LB}}(X)^{2}\right].$
@@ III-B2 Diffusion approximation
Let $\{W_{t}\}_{t\geq 0}$ be standard Brownian motion. By Donsker’s invariance principle [Donsker1951invariance], the rescaled battery process converges in distribution as $T\to\infty$:
@@
$B_{t}^{\mathrm{my}}\approx\mu_{B_{\mathrm{my}}}\,t+\sigma_{B_{\mathrm{my}}}\,W_{t}\sim\mathcal{N}\!\left(\mu_{B_{\mathrm{my}}}\,t,\;\sigma_{B_{\mathrm{my}}}^{2}\,t\right).$
Then, we show in Lem. 3 and (75) in the Appendix, the deficit $D_{T}^{\mathrm{my}}$ has a closed-form cumulative distribution function and expectation on its support $z\geq 0$.
@@
Thus system designs may typically be tuned near balance where $\overline{C}_{\mathrm{LB}}\approx\overline{R}$, the running minimum is fluctuation-dominated. In particular, at zero drift,
@@ III-C Routing error as a function of fluctuations
This comparison clarifies the relative importance of routing accuracy and robustness. When the cumulative routing error $\bar{K}\,\mathbb{E}_{\pi,X}[\Delta E_{i}^{\pi}(X)]\,T$ is significant relative to the fluctuation scale $\sigma_{B_{\mathrm{my}}}\sqrt{T}$, first-order drift dominates and improving routing accuracy yields the largest gains. When the two terms are comparable, fluctuation effects contribute non-negligibly to the reserve and policies may focus on accounting for variance in addition to mean optimality.
@@
Significant differences in relative LRM energy efficiency on tasks motivate highly accurate and efficient task dispatching. However, the dispatcher itself consumes energy and time. Let $E_{\pi}^{\mathrm{self}}(X)$ and $\xi_{\pi}^{\mathrm{self}}(X)$ denote the energy and latency to process task $X$ and select a model, under policy $\pi$. The dispatcher latency reduces available slack, $\sigma(X,t)\to\sigma(X,t)-\xi_{\pi}^{\mathrm{self}}(X)$, potentially shrinking the feasible set $\mathcal{M}_{F}(X,t)$ and precluding lower-energy allocations. The total energy overhead per task is then $E_{\pi}^{\mathrm{self}}(X)+\Delta E_{i}^{\pi}(X)$: the cost of routing plus the excess from any suboptimal selection. A more capable router may reduce $\operatorname{\mathbb{E}}[\Delta E_{i}^{\pi}(X)]$ but at the expense of increased $\operatorname{\mathbb{E}}[E_{\pi}^{\mathrm{self}}(X)]$ and $\operatorname{\mathbb{E}}[\xi_{\pi}^{\mathrm{self}}(X)]$. This tradeoff motivates the use of explicit and accurate scaling laws to simplify the dispatch, by reducing the mean problem for the dispatcher to difficulty prediction, enabling lightweight routing without heavy precomputation at dispatch time.
@@ IV Compute Scaling
In the preceding analysis, we treated the per-task success behavior, completion time, and energy expenditure as known to the dispatcher. In practice, these quantities can be estimated using predictors that encode task difficulty and model response. To ground this assumption, we adopt empirical scaling laws for large transformer architectures [Kaplan2020, Hoffmann2022] together with a recent theoretical extension for reasoning-style inference [EllisMohr2026TheoryInferenceCompute]. This makes the dependence of feasibility and energy on task requirements explicit.
@@ IV-A Token-based energy and latency
When a task $X$ is dispatched to model $M_{i}$, the router allocates either a thinking time $\tau$ or, equivalently, an output-token budget $\Omega\in\mathbb{N}$. For analysis, we treat $\Omega$ as a continuous proxy and use $\mathfrak{(\cdot)}$ to denote the resulting continuous-time (or continuous-token) quantities; the discrete-time quantities used elsewhere in the paper are recovered by sampling at a chosen resolution.
@@
$\mathfrak{e}_{i,L_{\mathrm{ctx}}}$ $=\alpha_{i}+2\beta_{i}\,L_{\mathrm{ctx}},$
$\xi_{i,L_{\mathrm{ctx}}}$ $=a_{i}+2b_{i}\,L_{\mathrm{ctx}},$
which correspond to parameter loading and feedforward/projection work (the constant terms) and attention over cached key-value pairs (the terms proportional to $L_{\mathrm{ctx}}$). Neglecting initial context for simplicity, we take $L_{\mathrm{ctx}}$ to grow proportionally with the number of generated tokens. Summing over $\Omega$ tokens yields the total energy and time:
$\mathfrak{E}_{i}(\Omega)$ $\coloneqq\sum_{v=0}^{\Omega-1}\mathfrak{e}_{i,\,v}\approx\alpha_{i}\,\Omega+\beta_{i}\,\Omega^{2},$
@@ IV-B Discretization
To obtain the discrete per-step energy profile $u\mapsto e_{i}(X,u)$ for our battery dynamics, define an auxiliary continuous time variable $\mathfrak{u}\in[0,\mathfrak{T}_{i}^{*}(X)]$. Since $\mathfrak{T}_{i}(\Omega)$ is strictly increasing for $\Omega\geq 0$, it admits an inverse. Letting $\Omega_{i}(\mathfrak{u})\coloneqq\mathfrak{T}_{i}^{-1}(\mathfrak{u})$ and $\mathfrak{E}_{i}(\mathfrak{u})\coloneqq\mathfrak{E}_{i}(\Omega_{i}(\mathfrak{u}))$, we define for $u=1,\dots,\tau_{i}^{*}(X),$
$e_{i}(X,u)\coloneqq\mathfrak{E}_{i}(u\Delta)-\mathfrak{E}_{i}((u-1)\Delta).$
By construction, $E_{i}^{*}(X)=\sum_{u=1}^{\tau_{i}^{*}(X)}e_{i}(X,u)$, allowing us to analyze routing for differing model sizes and capabilities under our mathematical framework.
@@ IV-C Scaling laws
To provide an explicit, model- and task-dependent expenditure shape induced by model size and inference scaling, we use the Directed Stochastic Skill Search inference scaling framework from [EllisMohr2026TheoryInferenceCompute]. Following that work, we may regard the task descriptor as $\theta=(l,m)$, where $m\in\mathbb{N}$ is the number of sequential skills and $l\in\mathbb{R}_{+}$ parameterizes their difficulty. Then for chain-of-thought reasoning using $\Omega_{i}^{*}(X)$ tokens,
$\psi_{i}(X,\Omega_{i}^{*}(X))=I_{\mathfrak{p}_{i}(l)}(m,\Omega_{i}^{*}/\omega-m+1),$
@@
To realize $\mathfrak{p}_{i}$ as a function of model size, we adopt training-compute optimal scaling as discussed by Hoffman et al. [Hoffmann2022]. Selecting dataset size along the training-compute-optimal frontier, the token-level pretraining loss may be written as $\mathcal{L}_{i}(N_{i})=\mathcal{L}_{\mathrm{irr}}+\Gamma N_{i}^{-\gamma}$ where $\mathcal{L}_{\mathrm{irr}}$ is an irreducible loss and $\Gamma,\gamma$ are based on model family efficiency for the given dataset. Then one simple functional form to model capability for a model of size $N_{i}$ is a sigmoid such that more difficult tasks (lower $l$) have a lower probability of success and vice versa:
$\mathfrak{p}_{i}(l)=\mathrm{Sigmoid}\bigl(\mathfrak{b}(l-\mathcal{L}_{i}(N_{i})\bigr)),$
@@ V Discussion
Analyses in Sec. IIIIV provide first- and second-order characterizations for the energy-aware LRM routing problem and further incude LRM scaling laws (both training-compute and inference-compute). As shown, the specific selection of dispatcher policy is critical to minimizing auxiliary energy usage.
Future work may consider alternative dispatch policies. Note the formal similarities between the present problem and joint routing-scheduling in energy-harvesting communication networks [Calvo-FullanaAMR2018], for which backpressure-type algorithms [TassiulasE1990] are developed. In our setting, a backpressure policy would route tasks by treating the request queue and the battery’s energy deficit as competing pressures. This algorithm would aggressively use large, energy-intensive models to clear backlogs when renewable energy is abundant, while using efficient models during energy shortages.
The computation graphs of common inference-compute scaling techniques are tree-structured, but there may be settings with more general directed acyclic graphs that impose more complicated precedence structures, cf. [LechowiczSBHWD2025]. Our current work treats LRM tasks as independent from one another, but future work may consider dependencies that add more constraints while simultaneously enabling new opportunities for energy optimization. Additional considerations include relaxing the i.i.d. assumptions on task arrivals and energy harvesting, which in practice may depend on temporal factors such as time of day and day of week, among other structural features. Studying limited-capacity or leaky battery models also presents further avenues for exploration.
Furthermore, we note that currently we treat the tolerance constraint $\varepsilon$ in (3) as deterministic. But, there may be settings where it may be stochastic as part of contracting or terms of use mechanisms for LRMs. Indeed, [Varaiya2011RiskLimitedDispatchEnergy, Varshney2014RiskLimitedDispatchKnowledge] introduce a probability distribution on risk tolerance and construct contract mechanisms with tranches of different risks.
@@ Appendix A Supporting Results and Proofs
$S_{t}\coloneqq\sum_{s=0}^{t-1}G_{s},\qquad t=0,1,\dots,T,$
@@
$D_{t+1}$ $=\bigl(-\mu_{t+1}\bigr)^{+}$ (40)
$=\max\Bigl\{0,\;-\min\{\mu_{t},B_{t+1}\}\Bigr\}$ (41)
@@
$\sum_{u=1}^{(t-s(x)+1)\wedge\tau(x)}\!\!\!\!e_{\mathrm{LB}}(x,u)\;$ $\leq\;\sum_{u=1}^{\tau(x)}e_{\mathrm{LB}}(x,u)$ (46)
@@
$\operatorname{Var}\left(\sum_{t=0}^{T-1}C_{t}^{\mathrm{MY}}\right)=\overline{K}T\cdot\operatorname{\mathbb{E}}\left[E_{\mathrm{LB}}(X)^{2}\right].$ (48)
@@
$\operatorname{Var}(S\mid N_{T}=n)$ $=\operatorname{Var}\left(\sum_{i=1}^{n}E_{\mathrm{LB}}(X_{i})\right)$ (51)
$=\sum_{i=1}^{n}\operatorname{Var}(E_{\mathrm{LB}}(X_{i}))$ (52)
$=n\cdot\operatorname{Var}(E_{\mathrm{LB}}(X)).$ (53)
@@
$\operatorname{Var}(S)$ $=\overline{K}T\left(\operatorname{Var}(E_{\mathrm{LB}}(X))+\left(\operatorname{\mathbb{E}}[E_{\mathrm{LB}}(X)]\right)^{2}\right)$ (55)
@@
$D_{T}\coloneqq\Bigl(-\min_{0\leq t\leq T}B_{t}\Bigr)^{+}.$
@@
$\mathbb{P}(D_{T}\leq z)=\Phi\!\left(\!\frac{z+\mu T}{\sigma\sqrt{T}}\right)-e^{-2\mu z/\sigma^{2}}\!\Phi\!\left(\!\frac{\mu T-z}{\sigma\sqrt{T}}\right),$ (57)
@@
$=\mathbb{P}\!\left(-m_{T}\leq z\right)$ (59)
$=\mathbb{P}\!\left(m_{T}\geq-z\right).$ (60)
The $(\cdot)^{+}$ is redundant for $z\geq 0$.
Define the rescaled process $Y_{t}\coloneqq B_{t}/\sigma=W_{t}+\tilde{\mu}t$, where $\tilde{\mu}\coloneqq\mu/\sigma$. Then $m_{T}=\sigma\min_{0\leq t\leq T}Y_{t}$, and for $z\geq 0$,
$\mathbb{P}(m_{T}\geq-z)$ $=\mathbb{P}\!\left(\inf_{0\leq t\leq T}Y_{t}\geq-\frac{z}{\sigma}\right)$ (61)
$=1-\mathbb{P}\!\left(\inf_{0\leq t\leq T}Y_{t}\leq-\frac{z}{\sigma}\right),$ (62)
@@
The process $Y_{t}=W_{t}+\tilde{\mu}t$ is Brownian motion with drift $\tilde{\mu}$ and unit volatility, denoted $W^{(\tilde{\mu})}_{t}$ by Borodin and Salminen [Borodin2002Handbook]. Indeed by [Borodin2002Handbook, formula 1.2.4 (p. 257)], for drift parameter $\alpha\in\mathbb{R}$ and level $y\leq x$,
$\mathbb{P}_{x}\!\left(\inf_{0\leq s\leq t}W^{(\alpha)}_{s}\leq y\right)$ $=\tfrac{1}{2}\operatorname{Erfc}\!\left(\tfrac{x-y+\alpha t}{\sqrt{2t}}\right)$
@@
$\mathbb{P}\!\left(\inf_{0\leq s\leq T}Y_{s}\leq-\frac{z}{\sigma}\right)$ $=\frac{1}{2}(\operatorname{Erfc}\!\left(\frac{z+\mu T}{\sigma\sqrt{2T}}\right)$
$+e^{-2\mu z/\sigma^{2}}\operatorname{Erfc}\!\left(\frac{z-\mu T}{\sigma\sqrt{2T}}\right)).$ (64)
@@
$\frac{1}{2}\operatorname{Erfc}\!\left(\frac{u}{\sqrt{2}}\right)=1-\Phi(u),$
@@
$\mathbb{P}\!\left(\inf_{0\leq s\leq T}Y_{s}\leq-\frac{z}{\sigma}\right)$
@@
$\begin{split}&=\Phi\!\left(\tfrac{-z-\mu T}{\sigma\sqrt{T}}\right)\\ &\quad+e^{-2\mu z/\sigma^{2}}\Phi\!\left(\tfrac{\mu T-z}{\sigma\sqrt{T}}\right).\end{split}$ (65)
@@
$\mathbb{P}(D_{T}\leq z)$ $=1-\mathbb{P}\!\left(\inf_{0\leq s\leq T}Y_{s}\leq-\frac{z}{\sigma}\right)$
@@
$\mathbb{E}[D_{T}]=\sigma\sqrt{T}\int_{0}^{\infty}\Bigl[Q(a+b)+e^{-2ab}\Phi(a-b)\Bigr]\,db,$
@@
where $\phi(u)\coloneqq(2\pi)^{-1/2}e^{-u^{2}/2}$. Substituting (71) into (69) yields, for $a=\frac{\mu\sqrt{T}}{\sigma}\neq 0$,
@@
$\mathbb{E}[D_{T}]=$ $\sigma\sqrt{T}\,\phi\!\left(\frac{\mu\sqrt{T}}{\sigma}\right)$ (73)
- $\mu T\!\left(1-\Phi\!\left(\frac{\mu\sqrt{T}}{\sigma}\right)\right)$ (74)
$+$ $\frac{\sigma^{2}}{2\mu}\!\left(2\Phi\!\left(\frac{\mu\sqrt{T}}{\sigma}\right)-1\right).$ (75)
@@
$\mathbb{E}[D_{T}]=-\mu T-\frac{\sigma^{2}}{2\mu}=|\mu|T+\frac{\sigma^{2}}{2|\mu|}.$
@@
$\mathbb{E}[D_{T}]\to\sigma\sqrt{T}\bigl(\phi(0)+\phi(0)\bigr)=\sigma\sqrt{\frac{2T}{\pi}},$

### v2

@@ III-B1 Battery random walk
$B_{t}^{\mathrm{my}}\;=\;\sum_{u=0}^{t-1}\bigl(R_{u}-C_{u}^{\mathrm{my}}\bigr)$ (26)
@@
$=t\,\overline{K}\,\operatorname{\mathbb{E}}_{X}\!\left[E_{\mathrm{LB}}(X)^{2}\right].$ (31)
Thus, the lumped myopic battery process is described by the mean, $\mu_{B_{\mathrm{my}}}$, and variance, $\sigma_{B_{\mathrm{my}}}^{2}$:
$\mu_{B_{\mathrm{my}}}\coloneqq\tfrac{1}{t}\operatorname{\mathbb{E}}[B_{t}^{\mathrm{my}}]=\overline{R}-\overline{C}_{\mathrm{LB}}$ (32)
$\sigma_{B_{\mathrm{my}}}^{2}\coloneqq\tfrac{1}{t}\operatorname{Var}(B_{t}^{\mathrm{my}})=\sigma_{R}^{2}+\overline{K}\operatorname{\mathbb{E}}_{X}\!\left[E_{\mathrm{LB}}(X)^{2}\right].$ (33)
@@ III-B2 Diffusion approximation
Let $\{W_{t}\}_{t\geq 0}$ be standard Brownian motion. By Donsker’s invariance principle [8], the rescaled battery process converges in distribution as $T\to\infty$:
@@
$B_{t}^{\mathrm{my}}\approx\mu_{B_{\mathrm{my}}}\,t+\sigma_{B_{\mathrm{my}}}\,W_{t}\sim\mathcal{N}\bigl(\mu_{B_{\mathrm{my}}}\,t,\;\sigma_{B_{\mathrm{my}}}^{2}\,t\bigr).$ (36)
Then, we show in Lem. 3 and (91), in the Appendix, the deficit $D_{T}^{\mathrm{my}}$ has a closed-form cumulative distribution function and expectation on its support $z\geq 0$.
@@
Thus, system designs may typically be tuned near balance where $\overline{C}_{\mathrm{LB}}\approx\overline{R}$, the running minimum is fluctuation-dominated. In particular, at zero drift,
@@ III-C Routing error as a function of fluctuations
This comparison clarifies the relative importance of routing accuracy and robustness. When the cumulative routing error $\overline{K}\,\mathbb{E}_{\pi,X}[\Delta E_{i}^{\pi}(X)]\,T$ is significant relative to the fluctuation scale $\sigma_{B_{\mathrm{my}}}\sqrt{T}$, first-order drift dominates and improving routing accuracy yields the largest gains. When the two terms are comparable, fluctuation effects contribute non-negligibly to the reserve and policies may focus on accounting for variance in addition to mean optimality.
@@
Significant differences in relative LRM energy efficiency on tasks motivate highly accurate and efficient task dispatching. However, the dispatcher itself consumes energy and time. Let $E_{\pi}^{\mathrm{self}}(X)$ and $\xi_{\pi}^{\mathrm{self}}(X)$ denote the energy and latency to process task $X$ and select a model, under policy $\pi$. The dispatcher latency reduces available slack, $\sigma(X,t)\to\sigma(X,t)-\xi_{\pi}^{\mathrm{self}}(X)$, potentially shrinking the feasible set $\mathcal{M}_{F}(X,t)$ and precluding lower-energy allocations. The total energy overhead per task is then $E_{\pi}^{\mathrm{self}}(X)+\Delta E_{i}^{\pi}(X)$: the cost of routing plus the excess from any suboptimal selection. A more capable router may reduce $\operatorname{\mathbb{E}}[\Delta E_{i}^{\pi}(X)]$ but at the expense of increased $\operatorname{\mathbb{E}}[E_{\pi}^{\mathrm{self}}(X)]$ and $\operatorname{\mathbb{E}}[\xi_{\pi}^{\mathrm{self}}(X)]$. While this overhead is likely negligible relative to the difference between executing large models, it may become non-negligible if improved routing accuracy requires a higher-capacity learned router. This tradeoff motivates the use of explicit and accurate scaling laws to simplify the dispatch, reducing the dispatch-time burden while preserving accuracy.
@@ IV Compute Scaling
In the preceding analysis, we treated the per-task success behavior, completion time, and energy expenditure as known to the dispatcher. In practice, these quantities can be estimated using predictors that encode task difficulty and model response. To ground this assumption, we adopt empirical scaling laws for large transformer architectures [15, 11] together with a recent theoretical extension for reasoning-style inference [9]. In turn, the dispatch variables $(s(x),i(x),\tau(x))$ map these task- and model-level predictors into system-level completion and consumption dynamics (e.g., $C_{t}$ and the induced battery drift/variability) discussed under the preceding regime analysis. This makes the dependence of feasibility and energy on task requirements explicit.
@@ IV-A Token-based energy and latency
When a task $X$ is dispatched to model $M_{i}$, the router allocates either a thinking time $\tau$ or, equivalently, an output-token budget $\Omega\in\mathbb{N}$. For analysis, we treat $\Omega$ as a continuous proxy and use fraktor symbols to denote the resulting continuous-time (or continuous-token) quantities; the discrete-time quantities used elsewhere in the paper are recovered by sampling at a chosen resolution.
@@
$\mathfrak{e}_{i,L_{\mathrm{ctx}}}=\alpha_{i}+2\beta_{i}\,L_{\mathrm{ctx}},\quad\xi_{i,L_{\mathrm{ctx}}}=a_{i}+2b_{i}\,L_{\mathrm{ctx}},$ (40)
which correspond to parameter loading and feedforward/projection work (the constant terms) and attention over cached key-value pairs (the terms proportional to $L_{\mathrm{ctx}}$). While inference includes a prefill cost determined by the initial prompt, we omit this term for clarity and because our focus is on reasoning workloads where decode-time compute dominates and scales with the number of generated tokens. Even when decode is termwise linear and often memory-bandwidth bound due to streaming model weights and KV-cache reads, the cached context grows over long reasoning traces, so summing these per-token memory and compute costs over generated tokens produces an effective quadratic dependence on the total generation length. Then, we take $L_{\mathrm{ctx}}$ to grow proportionally with the number of generated tokens. Summing over $\Omega$ tokens yields the total energy and time:
$\mathfrak{E}_{i}(\Omega)$ $\coloneqq\sum_{v=0}^{\Omega-1}\mathfrak{e}_{i,\,v}\approx\alpha_{i}\,\Omega+\beta_{i}\,\Omega^{2},$ (41)
@@ IV-B Discretization
To obtain the discrete per-step energy profile $u\mapsto e_{i}(X,u)$ for our battery dynamics, define an auxiliary continuous time variable $\mathfrak{u}\in[0,\mathfrak{T}_{i}^{*}(X)]$. Since $\mathfrak{T}_{i}(\Omega)$ is strictly increasing for $\Omega\geq 0$, it admits an inverse. Letting $\Omega_{i}(\mathfrak{u})\coloneqq\mathfrak{T}_{i}^{-1}(\mathfrak{u})$ and $\mathfrak{E}_{i}(\mathfrak{u})\coloneqq\mathfrak{E}_{i}(\Omega_{i}(\mathfrak{u}))$, we define for $u=\{0,\dots,\tau_{i}^{*}(X)-1\},$
$e_{i}(X,u)\coloneqq\mathfrak{E}_{i}((u+1)\Delta)-\mathfrak{E}_{i}(u\Delta).$ (43)
By construction, $E_{i}^{*}(X)=\sum_{u=0}^{\tau_{i}^{*}(X)-1}e_{i}(X,u)$, allowing us to analyze routing for differing model sizes and capabilities under our mathematical framework.
@@ IV-C Scaling laws
To provide an explicit, task-dependent expenditure induced by model size and inference scaling, we use the Directed Stochastic Skill Search inference scaling framework from [9]. Following that work, we may regard the task descriptor as $\theta=(l,m)$, where $m\in\mathbb{N}$ is the number of sequential skills and $l\in\mathbb{R}_{+}$ parameterizes their difficulty. Then for chain-of-thought reasoning using $\Omega_{i}^{*}(X)$ tokens,
$\psi_{i}(X;\Omega_{i}^{*}(X))=I_{\mathfrak{p}_{i}(l)}(m,\Omega_{i}^{*}/\omega-m+1),$ (44)
@@
To realize $\mathfrak{p}_{i}$ as a function of model size, we adopt training-compute optimal scaling as discussed by Hoffmann et al. [11]. Selecting dataset size along the training-compute-optimal frontier, the token-level pretraining loss may be written as $\mathcal{L}_{i}(N_{i})=\mathcal{L}_{\mathrm{irr}}+\Gamma N_{i}^{-\gamma}$ where $\mathcal{L}_{\mathrm{irr}}$ is an irreducible loss and $\Gamma,\gamma$ are based on model family efficiency for the given dataset. Then one simple functional form to model capability for a model of size $N_{i}$ is a sigmoid such that more difficult tasks (lower $l$) have a lower probability of success and vice versa:
$\mathfrak{p}_{i}(l)=\mathrm{Sigmoid}\bigl(\mathfrak{b}(l-\mathcal{L}_{i}(N_{i}))\bigr),$ (45)
@@ V Discussion
Analyses in Sec. IIIIV provide first- and second-order characterizations for the energy-aware LRM routing problem and further include LRM scaling laws (both training-compute and inference-compute). As shown, the specific selection of dispatcher policy is critical to minimizing auxiliary energy usage.
Future work may consider alternative dispatch policies. Note the formal similarities between the present problem and joint routing-scheduling in energy-harvesting communication networks [5], for which backpressure-type algorithms [27] are developed. In our setting, a backpressure policy would route tasks by treating the request queue and the battery’s energy deficit as competing pressures. This algorithm would aggressively use large, energy-intensive models to clear backlogs when renewable energy is abundant, while using efficient models during energy shortages.
The computation graphs of common inference-compute scaling techniques are tree-structured, but there may be settings with more general directed acyclic graphs that impose more complicated precedence structures, cf. [17]. Our current work treats LRM tasks as independent from one another, but future work may consider dependencies that add more constraints while simultaneously enabling new opportunities for energy optimization. Renewable supply is often temporally structured and correlated (e.g., wind and solar), and prior work has emphasized the importance of accounting for such temporal correlation in sustainable system design [10]. Such correlation can shift reserve and storage requirements.
At the same time, many correlations are partly predictable (e.g., diurnal/seasonal structure and site-dependent means) and can be mitigated operationally by shifting flexible workloads across time and, when available, across geographically distributed sites [7]. After accounting for this predictable structure, we conjecture the remaining short-timescale mismatch is closer to the i.i.d. baseline used here. We use the i.i.d. model primarily to obtain closed-form scaling and a clear baseline that clarifies the drift–fluctuation terms. Incorporating correlated and location-dependent energy models is an important direction for future work. Studying limited-capacity or leaky battery models also presents further avenues for exploration.
Furthermore, we note that currently we treat the tolerance constraint $\varepsilon$ in (5) as deterministic. But, there may be settings where it may be stochastic as part of contracting or terms of use mechanisms for LRMs. Indeed, [30, 2] introduce a probability distribution on risk tolerance and construct contract mechanisms with tranches of different risks.
@@ Appendix A Supporting Results and Proofs
$S_{t}\coloneqq\sum\nolimits_{s=0}^{t-1}G_{s},\qquad t=0,1,\dots,T,$ (46)
@@
$D_{t+1}$ $=\bigl(-\mu_{t+1}\bigr)^{+}=\Bigl(-\min\{\mu_{t},B_{t+1}\}\Bigr)^{+}$ (57)
@@
$\sum_{u=0}^{(t-s(x))\wedge(\tau(x)-1)}\!\!\!\!e_{\mathrm{LB}}(x,u)\;$ $\leq\;\sum_{u=0}^{\tau(x)-1}e_{\mathrm{LB}}(x,u)$ (63)
@@
$\operatorname{Var}\Bigl(\sum\nolimits_{t=0}^{T-1}C_{t}^{\mathrm{MY}}\Bigr)=\overline{K}T\cdot\operatorname{\mathbb{E}}\left[E_{\mathrm{LB}}(X)^{2}\right].$ (65)
@@
$\operatorname{Var}$ $(S\mid N_{T}=n)=\operatorname{Var}\Bigl(\sum_{i=0}^{n-1}E_{\mathrm{LB}}(X_{i})\Bigr)$ (68)
$=\sum_{i=0}^{n-1}\operatorname{Var}(E_{\mathrm{LB}}(X_{i}))=n\cdot\operatorname{Var}(E_{\mathrm{LB}}(X)).$ (69)
@@
$\operatorname{Var}(S)$ $=\overline{K}T\bigl(\operatorname{Var}(E_{\mathrm{LB}}(X))+\left(\operatorname{\mathbb{E}}[E_{\mathrm{LB}}(X)]\right)^{2}\bigr)$ (71)
@@
$D_{T}\coloneqq\Bigl(-\min_{0\leq t\leq T}B_{t}\Bigr)^{+}.$ (73)
@@
$\mathbb{P}(D_{T}\leq z)=\Phi\!\left(\!\frac{z+\mu T}{\sigma\sqrt{T}}\right)-e^{\frac{-2\mu z}{\sigma^{2}}}\!\Phi\!\left(\!\frac{\mu T-z}{\sigma\sqrt{T}}\right),$ (74)
@@
$=\mathbb{P}$ $\left(-m_{T}\leq z\right)=\mathbb{P}\!\left(m_{T}\geq-z\right).$ (76)
The $(\cdot)^{+}$ is redundant for $z\geq 0$. Define the rescaled process $Y_{t}\coloneqq B_{t}/\sigma=W_{t}+\tilde{\mu}t$, where $\tilde{\mu}\coloneqq\mu/\sigma$. Then $m_{T}=\sigma\min_{0\leq t\leq T}Y_{t}$, and for $z\geq 0$,
$\mathbb{P}(m_{T}\geq-z)$ $=\mathbb{P}\Bigl(\inf_{0\leq t\leq T}Y_{t}\geq-\frac{z}{\sigma}\Bigr)$ (77)
$=1-\mathbb{P}\Bigl(\inf_{0\leq t\leq T}Y_{t}\leq-\frac{z}{\sigma}\Bigr),$ (78)
@@
The process $Y_{t}=W_{t}+\tilde{\mu}t$ is Brownian motion with drift $\tilde{\mu}$ and unit volatility, denoted $W^{(\tilde{\mu})}_{t}$ by Borodin and Salminen [4]. Indeed by [4, formula 1.2.4 (p. 257)], for drift parameter $\alpha\in\mathbb{R}$ and level $y\leq x$,
$\mathbb{P}_{x}\Bigl(\inf_{0\leq s\leq t}W^{(\alpha)}_{s}\leq y\Bigr)$ $=\tfrac{1}{2}\operatorname{Erfc}\!\left(\tfrac{x-y+\alpha t}{\sqrt{2t}}\right)$
@@
$\mathbb{P}\Bigl(\inf_{0\leq s\leq T}Y_{s}\leq$- $\frac{z}{\sigma}\Bigr)=\frac{1}{2}\operatorname{Erfc}\!\left(\frac{z+\mu T}{\sigma\sqrt{2T}}\right)$
$+\frac{1}{2}e^{-2\mu z/\sigma^{2}}\operatorname{Erfc}\!\left(\frac{z-\mu T}{\sigma\sqrt{2T}}\right).$ (80)
@@
$\frac{1}{2}\operatorname{Erfc}\!\left(\frac{u}{\sqrt{2}}\right)=1-\Phi(u),$ (81)
@@
$\mathbb{P}\Bigl(\inf_{0\leq s\leq T}Y_{s}\leq-\frac{z}{\sigma}\Bigr)$
@@
$\begin{split}&=\Phi\!\left(\tfrac{-z-\mu T}{\sigma\sqrt{T}}\right)+e^{-2\mu z/\sigma^{2}}\Phi\!\left(\tfrac{\mu T-z}{\sigma\sqrt{T}}\right).\end{split}$ (82)
@@
$\mathbb{P}(D_{T}\leq$ $z)=1-\mathbb{P}\Bigl(\inf_{0\leq s\leq T}Y_{s}\leq-\frac{z}{\sigma}\Bigr)$
@@
$\mathbb{E}[D_{T}]\!=\!\sigma\sqrt{T}\int_{0}^{\infty}\!(Q(a\!+\!b)\!+\!e^{-2ab}\Phi(a\!-\!b))\,db,$ (86)
@@
where $\phi(u)\coloneqq(2\pi)^{-1/2}e^{-u^{2}/2}$. Substituting (88) and (89) into (87) yields, for $a=\frac{\mu\sqrt{T}}{\sigma}\neq 0$,
@@
$\mathbb{E}[D_{T}]=$ $\sigma\sqrt{T}\,\phi\bigl(\mu\sqrt{T}/\sigma\bigr)$ (91)
- $\mu T\bigl(1-\Phi\bigl(\mu\sqrt{T}/\sigma\bigr)\bigr)$
$+$ $\frac{\sigma^{2}}{2\mu}\bigl(2\Phi\bigl(\mu\sqrt{T}/\sigma\bigr)-1\bigr).$
@@
$\mathbb{E}[D_{T}]=-\mu T-\frac{\sigma^{2}}{2\mu}=|\mu|T+\frac{\sigma^{2}}{2|\mu|}.$ (93)
@@
$\mathbb{E}[D_{T}]\to\sigma\sqrt{T}\bigl(\phi(0)+\phi(0)\bigr)=\sigma\sqrt{2T/\pi},$ (94)